OpenStudy (howard-wolowitz):

Ok I got 97 / 7 for this: but I dont get how to round that

OpenStudy (howard-wolowitz):

do i put into decimal first?

OpenStudy (howard-wolowitz):

Ok, I think its B

OpenStudy (howard-wolowitz):

Its B

Nnesha (nnesha):

can you please explain how did you get that ?? :-)

OpenStudy (howard-wolowitz):

ok this is what I did: I worked the problem and then I converted 97/7 into a decimal

Nnesha (nnesha):

great i want to see that "work" how did you get 97/7 ?

OpenStudy (howard-wolowitz):

ok here: Rewrite log(7z+3)=2 in exponential form using the definition of a logarithm. If x and b are positive real numbers and b≠1, then logb(x)=y is equivalent to by=x. (10)2=7z+3 Remove parentheses around 10. 102=7z+3 Raise 10 to the power of 2 to get 100. 100=7z+3 Since z is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation. 7z+3=100 Since 3 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 3 from both sides. 7z=−3+100 Add −3 and 100 to get 97. 7z=97

OpenStudy (howard-wolowitz):

whichI subbed z into x

OpenStudy (howard-wolowitz):

so its z=97 /7

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Howard-Wolowitz ok here: Rewrite log(7z+3)=2 in exponential form using the definition of a logarithm. If x and b are positive real numbers and b≠1, then logb(x)=y is equivalent to by=x. (10)2=7z+3 Remove parentheses around 10. 102=7z+3 Raise 10 to the power of 2 to get 100. 100=7z+3 Since z is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation. 7z+3=100 Since 3 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 3 from both sides. 7z=−3+100 Add −3 and 100 to get 97. 7z=97 \(\color{blue}{\text{End of Quote}}\) that looks great gO_Od job!

OpenStudy (howard-wolowitz):

and if you put that into decimal form its : 13.85714286 : which is B

OpenStudy (howard-wolowitz):

ok awesome and thanks

Nnesha (nnesha):

np and please if you post ur answer with an explanation that would be great :-) thanks

OpenStudy (anonymous):

His actual name is Gregory Notowitz and he is a 45 year old Civics teacher for FLVS

OpenStudy (howard-wolowitz):

maybe ;)

Nnesha (nnesha):

ohh that's great SIR !

Nnesha (nnesha):

but i don't think so...

OpenStudy (howard-wolowitz):

Well doesn't matter, because you can think what you'd like! Its America!